# The foot, the fan, and the cuprate phase diagram: Fermi-volume-changing quantum phase transitions

### v1

@@ Abstract
A Fermi liquid with a ‘large’ Fermi surface (FL) can have a quantum phase transition to a spin density wave state (SDW) with reconstructed ‘small’ Fermi pockets. Both FL and SDW phases obey the Luttinger constraints on the volume enclosed by the Fermi surfaces. The critical spin fluctuations lead to spin-singlet $d$-wave pairing, as observed in the cuprates. Studies of the influence of spatial disorder on the SDW-FL quantum phase transition predict an extended quantum-critical Griffiths-type phase at low temperatures on the large Fermi surface side. These computations agree with the ‘foot’ seen in transport, and recent low temperature neutron scattering observations on La2-xSrxCuO4.
However, this theory cannot explain the higher temperature pseudogap and the ‘fan’ of strange metal of the hole-doped cuprates. Here we need to consider underlying Fermi-volume-changing quantum phase transitions without symmetry breaking. Then the small Fermi surface phase does not obey the Luttinger constraint, and is instead a ‘fractionalized Fermi liquid’ (FL*). A theory of FL* in single band models describes photoemission observations in the pseudogap phase. The FL*-FL quantum phase transition, and the influence of spatial disorder, is described by an analysis inspired by the Sachdev-Ye-Kitaev model. This theory successfully describes linear-in-temperature resistivity, optical conductivity and thermopower observations.
The crossovers connecting these lower and higher temperature descriptions are also discussed.
@@ 1 Jan Zaanen and the cuprates
I first met Jan Zaanen in the early nineties, and had innumerable discussions with him on a topic he was passionate about: the phase diagram of the cuprate high temperature superconductors.
@@
Figure 2: Phase diagram of the hole-doped cuprate superconductors from Ref. [2].
(See also the perspective of Proust and Taillefer [3], which has also influenced the present article.) Along with the observed low temperature phases with charge order and $d$-wave superconducting order, the diagram also shows two novel metallic phases, the pseudogap metal and the strange metal. Jan also thought a great deal about these metallic phases. He coined the term ‘Planckian’ [4] to highlight the similarity between aspects of the strange metal and charged black holes.
From today’s perspective, the important holographic model of strange metals by Jan and collaborators [5] can be presented as follows: the dopant electrons (visible in the ‘Fermi arc’ spectrum of the pseudogap metal [6, 7, 8, 9]) scatter strongly off a background two-dimensional black brane, leading to strange metal behavior. Specifically, they employed a charged black brane whose near-horizon physics is realized by an AdS2 metric in 1+1 spacetime dimensions, with the spatial co-ordinate of AdS2 being the emergent direction orthogonal to the black brane. Along with Ref. [10], this was the first time I had heard of AdS2 and its remarkable properties. Soon after, I realized [11] that the properties of AdS2 were strongly reminiscent of a class of models [12] now known as Sachdev-Ye-Kitaev models. Specifically, the common features were [11]:
@@
    A 1+1 dimensional conformal structure of the Green’s functions [13]. This implied that a suitably defined relaxation time $\tau(\omega)$ obeyed Planckian scaling as a function of frequency, $\omega$, and temperature, $T$:
@@
Returning to applications of the SYK model to the cuprates, while the scaling result in (1) is precisely that observed in optical conductivity measurements in the cuprates [19] (see Fig. 11; see also Ref. [20]), the large entropies in (2,4,5) are far from any observations. Moreover, the AdS2 holographic models (and also theories of critical Fermi surfaces of electrons coupled to gapless bosons [21, 22]) have the artifact of a continuous translational symmetry, which leads to perfect electrical conductivity, rather than strange metal behavior [23]. Much effort has been expended in overcoming these difficulties in applying these ideas to the cuprates, proceeding along two distinct directions. One direction modifies the holographic black brane models with either a periodic or random potential [24, 25, 26, 27]. The other direction works directly with the SYK model, and modifies it to a more realistic representation of the underlying lattice scale physics, as reviewed in Ref. [16]. It is this direction which we will discuss in the rest of this article. A fruitful formulation, which has been the focus of recent work, has been the extension to the two-dimensional complex Yukawa-Sachdev-Ye-Kitaev (2dYSYK) model [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 21, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52], which we will discuss in Section 4. A possible holographic dual of the 2dYSYK model remains an interesting open question for future work, which we will not address here.
@@ 2 The ‘foot’ and the ‘fan’
Figure 3: Color density plot of the resistivity exponent $\rho(T)\sim T^{n}$ in La2-xSrxCuO4 from Cooper et al. [53]. A magnetic field has been applied to suppress the superconductivity. The ‘foot’ and ‘fan’ annotations have been added.
@@
Figure 4: Theories of the phase diagram of Fig. 2. The theories in Sections 3, 4, and 5 are labeled by the corresponding section numbers. The large Fermi surface photoemission plot is from Ref. [55], and the pseudogap small Fermi surfaces are from the theory of Ref. [56]
As displayed in Fig. 4, we explain these remarkable behaviors by quantum phase transitions (QPTs) involving Fermi volume change, one with symmetry breaking, and the other without. In Section 3 we discuss the low temperature ‘foot’ in terms of a conventional spin density wave (SDW) onset in a disordered metal [45, 50]. The SDW order is closely connected to the stripe order discussed by Jan. However the SDW order disappers at higher temperatures, and so cannot explain the small Fermi surfaces (or ‘Fermi arcs’) in the pseudogap regime. We claim that the appropriate description of the pseudogap is terms of a ‘fractionalized Fermi liquid’ (FL*) [57, 58] in which the Fermi surface does not enclose the Luttinger volume, a behavior enabled by a background of a spin liquid. This then sets up a crossover from a Fermi liquid (FL) to a FL* to explain the higher temperature ‘fan’; we describe this crossover in terms of an underlying FL-FL* QPT in Section 4. Additionally, as we lower the temperature from the FL* in the pseudogap, there should be a crossover to the conventional SDW or $d$-wave superconducting (d-SC) states at low $T$ (which are not described by the FL-FL* QPT theory): this requires confinement of the fractionalized excitations of FL*, and is discussed in Section 5.
@@ 3 The ‘foot’: QPT with symmetry breaking
For the QPT with symmetry breaking in Fig. 4, we consider the onset of conventional SDW order in a FL, as discussed originally by Hertz [59], but in the presence of spatial randomness.
@@
Here $\tau^{a}$ are the Pauli matrices, $s$ is the parameter employed to tune across the transition, and $g$ is the Yukawa coupling between the fermions and bosons. We have included two sources of spatial randomness (symbols in purple are fixed random functions of space, with no dynamics). The spatially random potential $v({\bm{r}})$, with ensemble averages $\overline{v({\bm{r}})}=0$, $\overline{v({\bm{r}})v({\bm{r}^{\prime}})}=v^{2}\delta({\bm{r}}-{\bm{r}^{\prime}})$, acts on the fermion density, and plays a central role in the theory of disorder-induced electron localization [60]. Such fermion localization effects are also present here, but all indications are that such effects are not important for the cuprates. Instead, our focus will be on the more relevant ‘Harris disorder’, induced by spatial randomness in the position of the quantum critical point. This is represented by the spatially random ‘mass’ $\delta s({\bm{r}})$ with $\overline{\delta s({\bm{r}})}=0$, $\overline{\delta s({\bm{r}})\delta s({\bm{r}^{\prime}})}=\delta s^{2}\delta({\bm{r}}-{\bm{r}^{\prime}})$ which can lead to localization of the boson $\phi$.
A model closely related to (7) has recently been studied by large scale, high precision quantum Monte Carlo simulations by Patel et al. [50]. Their results are largely consistent with the analysis of Ref. [45] which applies a Hartree-Fock approximation to the $\phi^{4}$ term, but treats disorder numerically exactly. We describe the latter analysis here. We integrate out the fermions (assuming fermionic eigenmodes remain extended), and considering the Landau-damped Hertz theory for the boson alone, in the presence of a random mass. We discretize on a lattice of sites (labeled by $j$), and write the SDW order parameters in terms of a real $\phi$ with a single index $a=1\ldots M$ with $M=12$. In this manner, we obtain the action
@@
where $\Omega$ is a Matsubara frequency at a temperature $T$. For simplicity, we have assumed a global O($M$) symmetry, but this assumption can relaxed without significantly modifying the results.
@@
Figure 5: Phase diagram of the hole-doped cuprates from Ref. [65], based on their neutron scattering observations.
The results [61, 62, 63, 64, 45, 50] show a $T=0$ QPT at some $s=s_{c}$, accompanied by an extended gapless quantum Griffiths phase for $s>s_{c}$, with physics similar to the random Ising model in a transverse field [66]. Fermion spectral functions and transport properties have been computed, and are qualitatively similar to the ‘foot’ of the strange metal in Fig. 3. Radaelli et al. [65] have recently observed an extended regime of critical spin fluctuations in La2-xSrxCuO4 (extending earlier observations [67, 68, 69]; see Fig. 5), and their results are in general agreement with analytic continuation in frequency of the results of Refs. [45, 50].
@@
We conclude this discussion of the SDW QPT in a disordered metal by highlighting the structure of the bosonic eigenmodes $\psi_{\alpha}$ near the quantum critical point.
Figure 6: Adapted from Ref. [45]. (a) Localization length $\mathcal{L}_{\alpha}$ of overdamped bosonic eigenmodes of $\widetilde{\mathcal{S}}_{\phi}$ in (9) as a function of their energy $e_{\alpha}$. (b,c,d) Pictures of the corresponding bosonic eignfunctions. The ‘foot’ is described by the localized bosons. The universal theory of the ‘fan’ in Refs. [39, 47] using the 2dYSYK model in (12) applies to the regime of extended bosons.
Typical results from Ref. [45] are shown in Fig. 6. At higher $e_{\alpha}$, we obtain extended bosonic eigenmodes, where disorder can be treated in a self-averaging manner. However, we observe a remarkable non-monotonic behavior in the localization length $\mathcal{L}_{\alpha}$ of the eigenmodes as $e_{\alpha}$ is reduced. The bosonic modes localize at an intermediate energy, but then the localization length increases logarithmically upon further lowering the energy. This logarithmic increase is characteristic of infinite-randomness fixed points [66, 63], and arises from the interaction of the incipient localized modes. It is these localized modes of SDW spin fluctuations in Fig. 6 which are responsible for the extended critical quantum Griffiths phase, and they lead to the ‘foot’ in Fig. 3, and the critical spin fluctuations of Fig. 5 [45, 50]. We note that these localized SDW modes have some resemblance to models of the strange metal using two-level systems [76, 77].
It would clearly be of interest in future experiments to obtain spatially resolved spin fluctuation spectra, to determine if the critical spin fluctuations observed in neutron scattering [65] have the spatial structure illustrated in Fig. 6. Inhomogeneity in the superconductivity has already been observed in scanning tunneling microscopy experiments [78, 79] in work with Jan playing a significant role.

### v5

@@ Abstract
A Fermi liquid with a ‘large’ Fermi surface (FL) can have a quantum phase transition to a spin density wave state (SDW) with reconstructed ‘small’ Fermi pockets. Both FL and SDW phases obey the Luttinger constraints on the volume enclosed by the Fermi surfaces. Critical spin fluctuations lead to spin-singlet $d$-wave pairing, as observed in the cuprates. Studies of the influence of spatial disorder on the FL-SDW quantum phase transition predict an extended quantum-critical Griffiths-type phase at low temperatures on the large Fermi surface side. These computations agree with the ‘foot’ of strange metal transport, and recent low temperature neutron scattering observations on La2-xSrxCuO4.
However, this theory cannot explain the higher temperature pseudogap and the ‘fan’ of strange metal behavior of the hole-doped cuprates. Here we need to consider underlying Fermi-volume-changing quantum phase transitions without symmetry breaking. Then the small Fermi surface phase does not obey the Luttinger constraint, and the pseudogap metal is described by thermal fluctuations above a ‘fractionalized Fermi liquid’ (FL*) or a ‘holon metal’, with the descriptions related by a duality on a background spin liquid. The quantum critical fan is described using a field theory for an underlying FL-FL* quantum phase transition in the presence of spatial disorder. This field theory can be mapped to a form which can be analyzed using the methods of the Sachdev-Ye-Kitaev model. Such an analysis successfully models linear-in-temperature resistivity, optical conductivity and thermopower observations in the quantum critical fan.
The confinement crossover connecting these lower and higher temperature descriptions is also discussed.
arXiv:2501.16417
@@ 1 Jan Zaanen and the cuprates
I met Jan Zaanen in the early nineties, and had innumerable discussions with him on a topic he was passionate about: the phase diagram of the cuprate high temperature superconductors.
@@
Figure 2: Phase diagram of the hole-doped cuprate superconductors from Keimer et al. [2].
(See also the perspective of Proust and Taillefer [3], which has also influenced the present article.) Along with the observed low temperature phases with charge order and $d$-wave superconducting order, the diagram also shows two novel metallic phases, the pseudogap metal and the strange metal. Jan also thought a great deal about these metallic phases. He coined the term ‘Planckian’ [4] to highlight the similarity between aspects of strange metals and charged black holes [5, 6].
From today’s perspective, the important holographic model of strange metals by Jan and collaborators [7] can be presented as follows: the dopant electrons (visible in the ‘Fermi arc’ spectrum of the pseudogap metal [8, 9, 10, 11]) scatter strongly off a background two-dimensional black brane, leading to strange metal behavior. Specifically, they employed a charged black brane whose near-horizon physics is realized by an AdS2 metric in 1+1 spacetime dimensions, with the spatial co-ordinate of AdS2 being the emergent direction orthogonal to the black brane. Along with Refs. [12, 13], this was the first time I had heard of AdS2 and its remarkable properties. Soon after, I realized [14] that the properties of AdS2 were strongly reminiscent of a class of models [15] now known as Sachdev-Ye-Kitaev models. Specifically, the common features were [14]:
@@
    A 1+1 dimensional conformal structure of the Green’s functions [16, 17]. This implied that a suitably defined relaxation time $\tau(\omega)$ obeyed Planckian scaling as a function of frequency, $\omega$, and temperature, $T$:
@@
Returning to applications of the SYK model to the cuprates, while the scaling result in (1) is precisely that observed in optical conductivity measurements in the cuprates [23] (see Fig. 13; see also Refs. [24, 25]), the large entropies in (2,4,5) are far from any observations. Moreover, the AdS2 holographic models (and also theories of critical Fermi surfaces of electrons coupled to gapless bosons [26, 27]) have the artifact of a continuous translational symmetry, which leads to perfect electrical conductivity, rather than strange metal behavior [6]. Much effort has been expended in overcoming these difficulties in applying these ideas to the cuprates, proceeding along two distinct directions.
One direction modifies the holographic black brane models with either a periodic or random potential [28, 29, 30, 31].
The other direction works directly with the SYK model, and modifies it to a more realistic representation of the underlying lattice scale physics, as reviewed in Ref. [20]. It is this direction which we will discuss in the rest of this article. A fruitful formulation, which has been the focus of recent work, has been the mapping of theories of quantum phase transitions in disordered metals to the two-dimensional complex Yukawa-Sachdev-Ye-Kitaev (2dYSYK) model [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 26, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57], which we will discuss in Section 4. With the benefit of hindsight, we can now see some of this progress could have been made after the early papers [15, 19, 16, 18] on what was then sometimes called the SY model. But the intervening developments on the connections between SYK, black holes, and holography were important in pointing to the correct direction, and Jan played a significant role in these. A possible holographic dual of the 2dYSYK model remains an interesting open question for future work, which we will not address here.
@@ 2 The ‘foot’ and the ‘fan’
Figure 3: Color density plot of the resistivity exponent $\rho(T)\sim T^{n}$ in La2-xSrxCuO4 from Cooper et al. [58]. A magnetic field has been applied to suppress the superconductivity, and the resistance has been extrapolated to zero field. The ‘foot’ and ‘fan’ annotations have been added.
@@
Figure 4: Theories of the phase diagram of Fig. 2. The theories in Sections 3, 4, and 5 are labeled by the corresponding section numbers. The large Fermi surface photoemission plot is from Platé et al. [60], and the pseudogap Fermi arcs are from Shen et al. [8]. The latter is described by dual theories of thermal fluctuations of a holon metal or FL*.
As displayed in Fig. 4, we explain these remarkable behaviors by quantum phase transitions (QPTs) involving Fermi volume change, one with symmetry breaking, and the other without.
Before diving into the details of the cuprate phase diagram, it is helpful consider simpler analogous phenomena which occur in the triangular lattice antiferromagnet KYbSe2 [61]. This compound has long-range magnetic order at low $T$, and associated spin waves are observed at low energies. However, at higher energies, the neutron scattering observations (see Fig. 5) “identify a diffuse continuum with a sharp lower bound within the measured spectra …The key features of the data are reproduced by Schwinger boson theory” [61] of fractionalized spinons.
Figure 5: From Scheie et al. [61]. Neutron scattering spectrum in the triangular lattice antiferromagnet KYbSe2 (top panel), compared with the Schwinger boson (SB) theory of Ghioldi et al. [62] and DMRG.
(See also observations in NaYbSe2 [63], which display a spinon continuum without long-range magnetic order, and are proposed to be a $\mathbb{Z}_{2}$ spin liquid [64, 65].)
We argue for a similar phenomenology for the cuprates in Fig. 4. At low temperatures and doping, we consider symmetry breaking with conventional spin and charge order. At higher temperatures and low doping, we consider fractionalized states which enable Fermi volume change without symmetry breaking.
In Section 3 we discuss the low temperature ‘foot’ in terms of a conventional spin density wave (SDW) onset in a disordered metal [49, 54]. The SDW order is closely connected to the stripe order discussed by Jan.
However the SDW order disappears at higher temperatures, and so cannot explain the small Fermi surfaces (or ‘Fermi arcs’) in the pseudogap regime. We describe this higher temperature pseudogap regime not in terms of one or more fluctuating conventional orders [66, 67], but as a non-zero temperature realization of a quantum phase without symmetry breaking which could be stable at $T=0$ under suitable conditions. In principle, multiple quantum phases could exist, each distinct at $T=0$, yet all capable of providing satisfactory theoretical descriptions of the higher temperature pseudogap metal when thermal and quantum fluctuations are taken into account. We consider theories of distinct $T=0$ states known as ‘fractionalized Fermi liquids’ (FL*) [68, 69] and ‘holon metals’ [70, 71, 72, 73, 74, 75] (see Fig. 14). Both the $T=0$ FL* and holon metal states have Fermi surfaces which do not enclose the Luttinger volume, and such behavior is only possible in the presence of a background spin liquid [68, 69, 76, 77]. The spin liquid takes dual forms in the two phases, as we will discuss in Section 5. We present an underlying FL-FL* QPT in Section 4, and argue that it provides a satisfactory description over the higher temperature quantum critical fan [56]. But we do not fully resolve the nature of the pseudogap metal between FL* and the holon metal.
Additionally, as we lower the temperature from the pseudogap, there is a crossover to the conventional SDW or $d$-wave superconducting (d-SC) states at low $T$ (which are not described by the FL-FL* QPT theory): this requires confinement of the fractionalized excitations of the spin liquid background of the pseudogap, and is discussed in Section 5.
@@ 3 The ‘foot’: QPT with symmetry breaking
For the QPT with symmetry breaking in Fig. 4, we consider the onset of conventional SDW order in a FL, as discussed originally by Hertz [78] and Millis [79], but in the presence of spatial randomness.
@@
Here $\tau^{a}$ are the Pauli matrices, $s$ is the parameter employed to tune across the transition, and $g$ is the Yukawa coupling between the fermions and bosons. We have included two sources of spatial randomness (symbols in purple are fixed random functions of space, with no dynamics). The spatially random potential $v({\bm{r}})$, with ensemble averages $\overline{v({\bm{r}})}=0$, $\overline{v({\bm{r}})v({\bm{r}^{\prime}})}=v^{2}\delta({\bm{r}}-{\bm{r}^{\prime}})$, acts on the fermion density, and plays a central role in the theory of disorder-induced electron localization [80]. Such fermion localization effects are also present here, but all indications are that such effects are not important for the cuprates. Instead, our focus will be on the more relevant ‘Harris disorder’, induced by spatial randomness in the position of the quantum critical point. This is represented by the spatially random ‘mass’ $\delta s({\bm{r}})$ with $\overline{\delta s({\bm{r}})}=0$, $\overline{\delta s({\bm{r}})\delta s({\bm{r}^{\prime}})}=\delta s^{2}\delta({\bm{r}}-{\bm{r}^{\prime}})$ which can lead to localization of the boson $\phi$, and so must be treated non-perturbatively.
Patel et al. [54] have recently been studied a model closely related to the 2dYSYK model (12) by large scale, high precision quantum Monte Carlo simulations (see also Ref. [53]). As we will discuss above (12), the random coupling model (12) is obtained after a mapping [43] from the random ‘mass’ model (7). Here we discuss the analysis of Ref. [49] which yields similar results: it applies a Hartree-Fock approximation to the $\phi^{4}$ term in the random mass model (7), but treats disorder numerically exactly. We integrate out the fermions (assuming fermionic eigenmodes remain extended), and considering the Landau-damped Hertz theory for the boson alone, in the presence of a random mass. We discretize on a lattice of sites (labeled by $j$), and write the SDW order parameters in terms of a real $\phi$ with a single index $a=1\ldots M$ with $M=12$. In this manner, we obtain the action
@@
where $\Omega$ is a Matsubara frequency at a temperature $T$, $\gamma$ is the Landau damping, and the $\Omega^{2}/c^{2}$ term has been inserted as a high frequency cutoff. For simplicity, we have assumed a global O($M$) symmetry, but this assumption can relaxed without significantly modifying the results.
@@
Figure 6: Phase diagram of the hole-doped cuprates from Radaelli et al. [85], based on their neutron scattering observations.
The results [81, 82, 83, 84, 49, 54] show a $T=0$ FL-SDW QPT at some $s=s_{c}$, accompanied by an extended gapless quantum Griffiths phase for $s>s_{c}$, with physics similar to the random Ising model in a transverse field [86]. Fermion spectral functions and transport properties have been computed, and are qualitatively similar to the ‘foot’ of the strange metal in Fig. 3, and consistent with the recent high field observations of Campbell et al. [87] relating strange metals to spin fluctuations. Tranquada et al. [88] have also connected strange metal behavior to inhomogeneous spin fluctuations. Radaelli et al. [85] have recently observed an extended regime of critical spin fluctuations in La2-xSrxCuO4 (extending earlier observations [89, 90, 91]; see Fig. 6), and their results are in general agreement with analytic continuation in frequency of the results of Refs. [49, 54]. Similar neutron scattering results appeared recently in Ref. [92].
@@
It would be interesting to study the origin of $d$-wave superconductivity from these disordered spin fluctuations, extending the theories of $d$-wave superconductivity from a clean Fermi liquid [99, 100].
We conclude this discussion of the FL-SDW QPT in a disordered metal by highlighting the structure of the bosonic eigenmodes $\psi_{\alpha}$ near the quantum critical point.
Figure 7: Adapted from Patel et al. [49]. (a) Localization length $\mathcal{L}_{\alpha}$ of overdamped bosonic eigenmodes of $\widetilde{\mathcal{S}}_{\phi}$ in (9) as a function of their energy $e_{\alpha}$. (b,c,d) Pictures of the corresponding bosonic eignfunctions. The ‘foot’ is described by the localized bosons. The universal theory of the ‘fan’ by Patel et al. [43] and Li et al. [51] employs the mapping to the 2dYSYK model in (12), and applies to the regime of extended bosons.
Typical results from Ref. [49] are shown in Fig. 7. At higher $e_{\alpha}$, we obtain extended bosonic eigenmodes, where disorder can be treated in a self-averaging manner after the mapping to the 2dYSYK model to be discussed near (12). However, we observe a remarkable non-monotonic behavior in the localization length $\mathcal{L}_{\alpha}$ of the eigenmodes as $e_{\alpha}$ is reduced, which requires a non-perturbative and non-self-averaging treatment of disorder in either the random mass form of (7), or the random coupling form of (12). The bosonic modes localize at an intermediate energy, but then the localization length increases logarithmically upon further lowering the energy. This logarithmic increase is characteristic of infinite-randomness fixed points [86, 83], and arises from the interaction of the incipient localized modes. It is these localized modes of SDW spin fluctuations in Fig. 7 which are responsible for the extended critical quantum Griffiths phase, and they lead to the ‘foot’ in Fig. 3, and the critical spin fluctuations of Fig. 6 [49, 54].
We emphasize that this localization of bosonic modes takes place while the fermionic eigenstates remain extended. This is similar to early work on the disordered Hubbard model in three dimensions [101]. The localized SDW modes also have some resemblance to models of the strange metal using two-level systems [102, 103, 104].
It would clearly be of interest in future experiments to obtain spatially resolved spin fluctuation spectra, and then determine if the critical spin fluctuations observed in neutron scattering [85] have the spatial structure illustrated in Fig. 7. Inhomogeneity in the superconductivity has already been observed in scanning tunneling microscopy experiments [105, 106] in work with Jan playing a significant role.